OpenStudy (anonymous):

Solve the equation 2 over 5 (x − 4) = 2x for x

OpenStudy (ash2326):

We have 2/5(x-4)=2x multiply both sides by 5 2(x-4)=10x we have 2x-8=10x Let's use the symmetric property now 10x=2x-8 subtract both sides by 2x 8x=-8 divide both sides by 8 we get x=-1

OpenStudy (anonymous):

2/5x-4=2xfindx 10x2-8x-2=o ab=20 b=8 f(10.-2) 10x2-10x)-(2x-2) 10x2-2andx-1 x=2/5andx=1